# Arithmetic aspects of discrete periodic Toda flows

### v2

@@ 2.2 Jacobi matrices
$\mathrm{Jac}_{n}^{\mathrm{per}}(R)=\left\{L_{t}(z)=\begin{bmatrix}I_{1}^{t}+V_{n}^{t}&1&&(-1)^{n-1}I_{n}^{t}V_{n}^{t}/z\\ I_{1}^{t}V_{1}^{t}&I_{2}^{t}+V_{1}^{t}&\ddots&\\ &\ddots&\ddots&1\\ (-1)^{n-1}z&&I_{n-1}^{t}V_{n-1}^{t}&I_{n}^{t}+V_{n-1}^{t}\end{bmatrix}\right\},$ (2.8)
@@
$M_{t}(z)=\begin{bmatrix}1&&&(-1)^{n-1}V_{n}^{t}/z\\ V_{1}^{t}&1&&\\ &\ddots&\ddots&\\ &&V_{n-1}^{t}&1\end{bmatrix}$ (2.10)
@@
$R_{t}(z)=\begin{bmatrix}I_{1}^{t}&1&&\\ &I_{2}^{t}&\ddots&\\ &&\ddots&1\\ \ (-1)^{n-1}z&&&I_{n}^{t}\end{bmatrix}.$ (2.11)
@@
${L}_{t+1}(z)=R_{t}(z)L_{t}(z)R_{t}^{-1}.$ (2.13)
It will be important to keep in mind the explicit shape of $L_{t}(z)$ and
${L}_{t+1}(z)=\begin{bmatrix}I_{1}^{t}+V_{1}^{t}&1&&(-1)^{n-1}I_{1}^{t}V_{n}^{t}/z\\ I_{2}^{t}V_{1}^{t}&I_{2}^{t}+V_{2}^{t}&\ddots&\\ &\ddots&\ddots&1\\ (-1)^{n-1}z&&I_{n}^{t}V_{n-1}^{t}&I_{n}^{t}+V_{n}^{t}\end{bmatrix}.$ (2.14)
@@
We introduce the notation
$\underline{L}_{t}(z)=L_{t+1}(z).$ (2.15)
Remark 2.2.
From now on we will (often) suppress the time parameter $t$. Unless stated otherwise, $I_{i}$ and $V_{j}$ always mean $I_{i}^{t}$ and $V_{j}^{t}$. We will also denote $L_{t}(z)$ and $\underline{L}_{t}(z)$ by $L$ and $\underline{L}$.
@@ 2.3 The spectral curve
$I_{o}=\prod_{i=1}^{n}I_{i}^{t}\ \ \ \text{ and }\ \ \ V_{o}=\prod_{i=1}^{n}V_{i}^{t}$ (2.17)
@@
$|\mathcal{L}|=(I_{n}+V_{n-1}+x)|\mathcal{L}_{1}|-I_{n-1}V_{n-1}|\mathcal{L}_{n-1,n}|+z|\mathcal{L}_{1,n}|.$ (2.18)
@@
$|\mathcal{L}_{n-1,n}|=|\mathcal{L}_{2}|-(I_{n}V_{n}/z)(I_{1}V_{1}\cdots I_{n-2}V_{n-2})$ (2.19)
@@
$h(x)=(I_{n}+V_{n-1}-x)|\mathcal{L}_{1}|-I_{n}V_{n}|^{1}\mathcal{L}_{1}|-I_{n-1}V_{n-1}|\mathcal{L}_{2}|\in R[x]$ (2.21)
@@
Define the polynomials
$u$ $=u^{t}(x)=|\mathcal{L}_{1}|\in R[x],$ (2.23)
$v$ $=v^{t}(x)=I^{t}_{n}V^{t}_{n}|{}^{1}\mathcal{L}_{1}|\in R[x],$ (2.24)
$w$ $=w^{t}(x)=I^{t}_{n-1}V^{t}_{n-1}|\mathcal{L}_{2}|\in R[x]$ (2.25)
and
$\underline{u}$ $=u^{t+1}(x)=|\underline{\mathcal{L}}_{1}|\in R[x],$ (2.26)
$\underline{v}$ $=v^{t+1}(x)=I^{t}_{1}V^{t}_{n}|{}^{1}\underline{\mathcal{L}}_{1}|\in R[x],$ (2.27)
$\underline{w}$ $=w^{t+1}(x)=I^{t}_{n}V^{t}_{n-1}|\underline{\mathcal{L}}_{2}|\in R[x].$ (2.28)
@@
$|{}^{k}\mathcal{L}_{l}|$ $=(I_{k+1}+V_{k}+x)|{}^{k+1}\mathcal{L}_{l}|-I_{k+1}V_{k+1}|{}^{k+2}\mathcal{L}_{l}|,$ (2.30)
$|{}^{k}\mathcal{L}_{l}|$ $=(I_{n-l}+V_{n-l-1}+x)|{}^{k}\mathcal{L}_{l+1}|-I_{n-l-1}V_{n-l-1}|{}^{k}\mathcal{L}_{l+2}|,$ (2.31)
|¯$|{}^{k}\underline{\mathcal{L}}_{l}|$ $=(I_{k+1}+V_{k+1}-x)|{}^{k+1}\underline{\mathcal{L}}_{l}|-I_{k+2}V_{k+1}|{}^{k+2}\underline{\mathcal{L}}_{l}|$ (2.32)
|¯$|{}^{k}\underline{\mathcal{L}}_{l}|$ $=(I_{n-l}+V_{n-l}-x)|{}^{k}\underline{\mathcal{L}}_{l+1}|-I_{n-l}V_{n-l-1}|{}^{k}\underline{\mathcal{L}}_{l+2}|.$ (2.33)
@@
2.3.1 Different models of $\mathcal{C}$
@@
2.3.2 Standard model $\widetilde{\mathcal{C}}$
@@
$g=\mathrm{deg}(h(x))-1=n-1.$ (2.36)
2.3.3 Imaginary model $\mathcal{C}_{\alpha}$
@@
$(x,y)\longmapsto(\tfrac{1}{x-\alpha},\tfrac{y}{\beta(x-\alpha)^{g+1}})=(x_{\alpha},y_{\alpha})$ (2.37)
@@
${\mathcal{C}}_{\alpha}:y_{\alpha}^{2}=F_{\alpha}(x_{\alpha})$ (2.38)
@@
2.4 The cyclic shift and some involutions
@@
$\sigma:(I_{i},V_{i})\longmapsto(I_{i+1},V_{i+1}),$ (2.39)
with the usual convention (2.4).
@@
The spectral curve $\mathcal{C}$ is invariant under the cyclic shift map $\sigma$.
@@
$h=\sigma(h)=(I_{1}+V_{n}+x)|\sigma(\mathcal{L})_{1}|-I_{1}V_{1}|{}^{1}\sigma(\mathcal{L})_{1}|-I_{n}V_{n}|\sigma(\mathcal{L})_{2}|.$ (2.40)
This follows from $|\sigma(\mathcal{L})_{1}|=|{{}^{1}\mathcal{L}|}$, $|{}^{1}\sigma(\mathcal{L})_{1}|=|{}^{2}\mathcal{L}|$ and $|\sigma(\mathcal{L})_{2}|=|{}^{1}\mathcal{L}_{1}|$ in combination with Lemma 2.1. ∎
@@
The map
${}^{\dagger}:(I_{1},\dots,I_{n},V_{1},\dots,V_{n})\mapsto(V_{n-1},\dots,V_{1},V_{n},I_{n-1},\dots,I_{1},I_{n})$ (2.41)
defines an involution on $R^{2n}$.
@@
The spectral curve is invariant under the involution .
@@
This follows from Lemma 2.4 and the last Lemma. ∎
@@
Define the polynomials
$\sigma(u)$ $=\sigma(u)^{t}(x)=|\sigma(\mathcal{L})_{1}|\in R[x],$ (2.42)
$\sigma(v)$ $=\sigma(v)^{t}(x)=I^{t}_{n}V^{t}_{n}|{}^{1}\sigma(\mathcal{L})_{1}|\in R[x],$ (2.43)
$\sigma(w)$ $=\sigma(w)^{t}(x)=I^{t}_{n-1}V^{t}_{n-1}|\sigma(\mathcal{L})_{2}|\in R[x]$ (2.44)
and
$u^{\dagger}$ $=u^{t,\dagger}(x)=|\mathcal{L}^{\dagger}_{1}|\in R[x],$ (2.45)
$v^{\dagger}$ $=v^{t,\dagger}(x)=I^{t}_{n}V^{t}_{n}|{}^{1}\mathcal{L}^{\dagger}_{1}|\in R[x],$ (2.46)
$w^{\dagger}$ $=w^{t,\dagger}(x)=I^{t}_{1}V^{t}_{1}|\mathcal{L}^{\dagger}_{2}|\in R[x].$ (2.47)
@@
$\sigma_{i,j}:I_{i}\leftrightarrow I_{j},V_{i}\leftrightarrow V_{j}.$ (2.48)
@@
$s_{i}:I_{j}\mapsto\left\{\begin{array}[]{c}-I_{i}\text{, if $j=i$}\\ \ I_{j}\text{, if $j\neq i$}\end{array}\right.$ (2.49)
@@
$t_{i}:V_{j}\mapsto\left\{\begin{array}[]{c}-V_{i}\text{, if $j=i$}\\ \ V_{j}\text{, if $j\neq i$}\end{array}.\right.$ (2.50)
2.5 Some useful observations
@@
$\begin{array}[]{ccc}u=\tfrac{h_{I_{n},V_{n-1}}}{x},&v=I_{n}V_{n}\tfrac{u_{I_{1},V_{n}}}{x},&w=I_{n-1}V_{n-1}\tfrac{u_{I_{n-1},V_{n-2}}}{x},\\ \\ \underline{u}=\tfrac{h_{I_{n},V_{n}}}{x},&\underline{v}=I_{1}V_{n}\tfrac{\underline{u}_{I_{1},V_{1}}}{x},&\underline{w}=I_{n}V_{n-1}\tfrac{\underline{u}_{I_{n-1},V_{n-1}}}{x},\\ \\ \sigma(u)=\tfrac{h_{I_{1},V_{n}}}{x},&\sigma(v)=I_{1}V_{1}\tfrac{\sigma(u)_{I_{2},V_{1}}}{x},&\sigma(w)=I_{n}V_{n}\tfrac{\sigma(u)_{I_{n},V_{n-1}}}{x},\\ \\ u^{\dagger}=\sigma(u),&v^{\dagger}=\sigma(w),&w^{\dagger}=\sigma(v).\end{array}$
@@
We will only prove the first two relation. The other relations follow similarly. The first follows trivially from the definition of $h$, see (2.29). The second one follows from the expansion
$|\mathcal{L}_{1}|=(I_{1}+V_{n}+x)|{}^{1}\mathcal{L}_{1}|-I_{1}V_{1}|{}^{2}\mathcal{L}_{1}|.$ (2.51)
@@
Setting $I_{i}$ equal to zero in the above Lemma can be expressed in terms of the involution $s_{i}$, namely we have
$u_{I_{i}}=\tfrac{u+s_{i}(u)}{2}.$ (2.53)
@@
$h,u,v,w,\underline{u},\underline{v},\underline{w},\sigma(u),\sigma(v),\sigma(w),u^{\dagger},v^{\dagger},w^{\dagger}\in R[x]$ (2.54)
@@
$h$ $=(I_{n}+V_{n-1}+x)u-v-w$ (2.55)
$=xu+I_{n}u_{V_{n}}+V_{n-1}u_{I_{n-1}}$ (2.56)
we only need to show that $u=|\mathcal{L}_{1}|$ is subtraction-free. This follows from the tridiagonal structure of $|\mathcal{L}_{1}|$ (and Lemma 2.1), as one can immediately see that the only negative terms appearing $|\mathcal{L}_{1}|$ are of the form $-I_{j}V_{j}|\mathcal{L}_{n-j+1}|$ and those are cancelled by the expressions $(I_{j}+V_{j-1}+x)(I_{j+1}+V_{j}+x)|\mathcal{L}_{n-j+1}|$. ∎
@@
The last Lemma is the main reason for our choice of normalization of $\mathcal{C}$, as subtraction-free polynomials can be nicely tropicalized. This will be used in section 3.6.
@@
$|{}^{1}\mathcal{L}_{1}||\mathcal{L}_{2}|-|\mathcal{L}_{1}||{}^{1}\mathcal{L}_{2}|=I_{1}V_{1}\cdots I_{n-2}V_{n-2}.$ (2.58)
@@
$|{}^{2}\mathcal{L}_{1}||{}^{1}\mathcal{L}_{2}|-|{}^{1}\mathcal{L}_{1}||{}^{2}\mathcal{L}_{2}|=I_{2}V_{2}\cdots I_{n-2}V_{n-2}.$ (2.59)
@@
$|{}^{n-2}\mathcal{L}_{1}||{}^{n-3}\mathcal{L}_{2}|-|{}^{n-3}\mathcal{L}_{1}||{}^{n-2}\mathcal{L}_{2}|=I_{n-2}V_{n-2}.$ (2.60)
@@
$I_{n}V_{n}|{}^{1}\mathcal{L}_{2}|=(I_{n}+V_{n-1}+x)\sigma^{-1}(v)-\sigma^{-1}(u)$ (2.61)
@@
This follows from $|{}^{1}\mathcal{L}_{2}|=|{}^{2}\sigma^{-1}(\mathcal{L})_{1}|$ and Lemma 2.1. ∎
@@
$I_{n}u-v$ $=I_{n}u_{V_{n}}$ (2.63)
$=I_{n}\underline{u}-\underline{w},$ (2.64)
$(V_{n-1}+x)u-w$ $=(V_{n}+x)\underline{u}-\underline{v}.$ (2.65)
@@
$u$ $=(I_{1}+V_{n}+x)|{}^{1}\mathcal{L}_{1}|-I_{1}V_{1}|{}^{2}\mathcal{L}_{1}|,$ (2.66)
$\underline{u}$ $=(I_{1}+V_{1}+x)|{}^{1}\underline{\mathcal{L}}_{1}|-I_{2}V_{1}|{}^{2}\underline{\mathcal{L}}_{1}|$ (2.67)
and Lemma 2.4. The second assertion follows from the first and $h=\underline{h}.$
@@
$u_{n}$ $=(u_{n})_{V_{n}}+V_{n}\sigma_{n-1}(u_{n-1}),$ (2.68)
$v_{n}$ $=I_{n}V_{n}\sigma_{n-1}(u_{n-1}),$ (2.69)
$w_{n}$ $=I_{n-1}V_{n-1}\sigma_{n-1,n}(u_{n-1}).$ (2.70)
@@
Let us start with the second relation. This follows from
$|{}^{1}\mathcal{L}_{1}|=|\sigma_{n}(\mathcal{L})_{2}|=|\sigma_{n-1}(\mathcal{L}_{n-1})_{1}|.$ (2.71)
@@
$u_{n}-(u_{n})_{V_{n}}=u_{n}-u_{n}+V_{n}|{}^{1}\mathcal{L}_{1}|$ (2.72)
and the second relation. The other relation follow similarly. ∎
@@
$(\underline{u}_{n})_{V_{n-1}}=(u_{n})_{V_{n}}.$ (2.73)
2.6 Mumford’s algebraic description of $\mathrm{Jac(\mathcal{C})}$

### v3

@@ 2.2. Jacobi matrices
$\mathrm{Jac}_{n}^{\mathrm{per}}(R)=\left\{L_{t}=\begin{bmatrix}I_{1}^{t}+V_{n}^{t}&1&&(-1)^{n-1}I_{n}^{t}V_{n}^{t}/z\\ I_{1}^{t}V_{1}^{t}&I_{2}^{t}+V_{1}^{t}&\ddots&\\ &\ddots&\ddots&1\\ (-1)^{n-1}z&&I_{n-1}^{t}V_{n-1}^{t}&I_{n}^{t}+V_{n-1}^{t}\end{bmatrix}\right\}$,
@@
$M_{t}=\begin{bmatrix}1&&&(-1)^{n-1}V_{n}^{t}/z\\ V_{1}^{t}&1&&\\ &\ddots&\ddots&\\ &&V_{n-1}^{t}&1\end{bmatrix}$
@@
$R_{t}=\begin{bmatrix}I_{1}^{t}&1&&\\ &I_{2}^{t}&\ddots&\\ &&\ddots&1\\ \ (-1)^{n-1}z&&&I_{n}^{t}\end{bmatrix}$.
@@
(2.3) ${L}_{t+1}=R_{t}L_{t}R_{t}^{-1}.$
It will be important to keep in mind the explicit shape of $L_{t}$ and
$\overline{L}_{t}={L}_{t+1}=\begin{bmatrix}I_{1}^{t}+V_{1}^{t}&1&&(-1)^{n-1}I_{1}^{t}V_{n}^{t}/z\\ I_{2}^{t}V_{1}^{t}&I_{2}^{t}+V_{2}^{t}&\ddots&\\ &\ddots&\ddots&1\\ (-1)^{n-1}z&&I_{n}^{t}V_{n-1}^{t}&I_{n}^{t}+V_{n}^{t}\end{bmatrix}$.
@@
From now on we will usually suppress the time parameter $t$. Unless stated otherwise, $I_{i}$ and $V_{j}$ always mean $I_{i}^{t}$ and $V_{j}^{t}$ and so on.
@@ 2.3. The spectral curve
$I_{\Pi}=\prod_{i=1}^{n}I_{i}^{t}\ \ \ \text{ and }\ \ \ V_{\Pi}=\prod_{i=1}^{n}V_{i}^{t}$
@@
$|\mathcal{L}|=(I_{n}+V_{n-1}+x)|\mathcal{L}_{1}|-I_{n-1}V_{n-1}|\mathcal{L}_{n-1,n}|+z|\mathcal{L}_{1,n}|.$
@@
$|\mathcal{L}_{n-1,n}|=|\mathcal{L}_{2}|-(I_{n}V_{n}/z)(I_{1}V_{1}\cdots I_{n-2}V_{n-2})$
@@
$h(x)=(I_{n}+V_{n-1}+x)|\mathcal{L}_{1}|-I_{n}V_{n}|^{1}\mathcal{L}_{1}|-I_{n-1}V_{n-1}|\mathcal{L}_{2}|\in R[x]$
@@
Define the polynomials $u,v,w,\overline{u},\overline{v},\overline{w}\in R[x]$ by
$u=|\mathcal{L}_{1}|,$ $v=I^{t}_{n}V^{t}_{n}|{}^{1}\mathcal{L}_{1}|,$ $w=I^{t}_{n-1}V^{t}_{n-1}|\mathcal{L}_{2}|$
$\overline{u}=|\overline{\mathcal{L}}_{1}|,$ $\overline{v}=I^{t}_{1}V^{t}_{n}|{}^{1}\overline{\mathcal{L}}_{1}|,$ $\overline{w}=I^{t}_{n}V^{t}_{n-1}|\overline{\mathcal{L}}_{2}|.$
@@
$|{}^{k}\mathcal{L}_{l}|$ $=(I_{k+1}+V_{k}+x)|{}^{k+1}\mathcal{L}_{l}|-I_{k+1}V_{k+1}|{}^{k+2}\mathcal{L}_{l}|,$
$|{}^{k}\mathcal{L}_{l}|$ $=(I_{n-l}+V_{n-l-1}+x)|{}^{k}\mathcal{L}_{l+1}|-I_{n-l-1}V_{n-l-1}|{}^{k}\mathcal{L}_{l+2}|,$
$|{}^{k}\overline{\mathcal{L}}_{l}|$ $=(I_{k+1}+V_{k+1}+x)|{}^{k+1}\overline{\mathcal{L}}_{l}|-I_{k+2}V_{k+1}|{}^{k+2}\overline{\mathcal{L}}_{l}|$
$|{}^{k}\overline{\mathcal{L}}_{l}|$ $=(I_{n-l}+V_{n-l}+x)|{}^{k}\overline{\mathcal{L}}_{l+1}|-I_{n-l}V_{n-l-1}|{}^{k}\overline{\mathcal{L}}_{l+2}|.$
@@
2.3.1. Different models of $\mathcal{C}$
@@
2.3.2. Standard model $\widetilde{\mathcal{C}}$
@@
$g=\mathrm{deg}(h(x))-1=n-1.$
2.3.3. Imaginary model $\mathcal{C}_{\alpha}$
@@
$(x,y)\longmapsto(\tfrac{1}{x-\alpha},\tfrac{y}{\beta(x-\alpha)^{g+1}})=(x_{\alpha},y_{\alpha})$
@@
${\mathcal{C}}_{\alpha}:y_{\alpha}^{2}=F_{\alpha}(x_{\alpha})$
@@
2.4. The cyclic shift and some involutions
@@
$\sigma:(I_{i},V_{i})\longmapsto(I_{i+1},V_{i+1}),$
with the usual convention (2.2). We write $(\mathbb{I},\mathbb{V})\mapsto(\mathbb{I}^{\sigma},\mathbb{V}^{\sigma})$.
@@
The spectral curve $\mathcal{C}$ is invariant under the cyclic shift map $\sigma$, i.e.,
$h^{\sigma}=h\text{ and }f^{\sigma}=f.$
@@
$h=h^{\sigma}=(I_{1}+V_{n}+x)|\mathcal{L}^{\sigma}_{1}|-I_{1}V_{1}|{}^{1}\mathcal{L}^{\sigma}_{1}|-I_{n}V_{n}|\mathcal{L}^{\sigma}_{2}|.$
This follows from $|\mathcal{L}^{\sigma}_{1}|=|{{}^{1}\mathcal{L}|}$, $|{}^{1}\mathcal{L}^{\sigma}_{1}|=|{}^{2}\mathcal{L}|$ and $|\mathcal{L}^{\sigma}_{2}|=|{}^{1}\mathcal{L}_{1}|$ in combination with Lemma 2.1. ∎
@@
The dagger involution $^{\dagger}:R^{2n}\longrightarrow R^{2n}$ is defined by
$^{\dagger}:(I_{1},\dots,I_{n},V_{1},\dots,V_{n})\mapsto(V_{n-1},\dots,V_{1},V_{n},I_{n-1},\dots,I_{1},I_{n}).$
We write $(\mathbb{I},\mathbb{V})\mapsto(\mathbb{I}^{\dagger},\mathbb{V}^{\dagger})$.
@@
The spectral curve is invariant under the dagger involution , i.e.,
h=h $^{\dagger}=h\text{ and }f^{\dagger}=f$.
@@
This follows from Lemma 2.4 in combination with Lemma 2.2
@@
Define the polynomials $u^{\sigma},v^{\sigma},w^{\sigma},u^{\dagger},v^{\dagger},w^{\dagger}\in R[x]$ by
$u^{\sigma}=|\mathcal{L}^{\sigma}_{1}|,$ $v^{\sigma}=I^{t}_{n}V^{t}_{n}|{}^{1}\mathcal{L}^{\sigma}_{1}|,$ $w^{\sigma}=I^{t}_{n-1}V^{t}_{n-1}|\mathcal{L}^{\sigma}_{2}|$
$u^{\dagger}=|\mathcal{L}^{\dagger}_{1}|,$ $v^{\dagger}=I^{t}_{n}V^{t}_{n}|{}^{1}\mathcal{L}^{\dagger}_{1}|,$ $w^{\dagger}=I^{t}_{1}V^{t}_{1}|\mathcal{L}^{\dagger}_{2}|.$
@@
$\sigma_{i,j}:I_{i}\leftrightarrow I_{j},V_{i}\leftrightarrow V_{j}.$
@@
$s_{i}:I_{j}\mapsto\left\{\begin{array}[]{c}-I_{i}\text{, if $j=i$}\\ \ I_{j}\text{, if $j\neq i$}\end{array}\right$.
@@
$t_{i}:V_{j}\mapsto\left\{\begin{array}[]{c}-V_{i}\text{, if $j=i$}\\ \ V_{j}\text{, if $j\neq i$}\end{array}.\right$.
2.5. Some useful observations
@@
$u$ $=\tfrac{h_{I_{n},V_{n-1}}}{x},$ $v$ $=I_{n}V_{n}\tfrac{u_{I_{1},V_{n}}}{x},$ $w$ $=I_{n-1}V_{n-1}\tfrac{u_{I_{n-1},V_{n-2}}}{x},$
$\overline{u}$ $=\tfrac{h_{I_{n},V_{n}}}{x},$ $\overline{v}$ $=I_{1}V_{n}\tfrac{\overline{u}_{I_{1},V_{1}}}{x},$ $\overline{w}$ $=I_{n}V_{n-1}\tfrac{\overline{u}_{I_{n-1},V_{n-1}}}{x},$
$u^{\sigma}$ $=\tfrac{h_{I_{1},V_{n}}}{x},$ $v^{\sigma}$ $=I_{1}V_{1}\tfrac{u^{\sigma}_{I_{2},V_{1}}}{x},$ $w^{\sigma}$ $=I_{n}V_{n}\tfrac{u^{\sigma}_{I_{n},V_{n-1}}}{x},$
$u^{\dagger}$ $=u^{\sigma},$ ${v^{\dagger}}$ $=w^{\sigma},$ $w^{\dagger}$ $=v^{\sigma}.$
@@
We will only prove the first two relations. The other relations follow similarly. The first follows trivially from the definition of $h$, see (2.4). The second one follows from the expansion
$|\mathcal{L}_{1}|=(I_{1}+V_{n}+x)|{}^{1}\mathcal{L}_{1}|-I_{1}V_{1}|{}^{2}\mathcal{L}_{1}|.$
@@
Setting $I_{i}=0$ above Lemma can be expressed in terms of the involution $s_{i}$, namely we have
$u_{I_{i}}=\tfrac{1}{2}(u+u^{s_{i}}).$
@@
$h,u,v,w,\overline{u},\overline{v},\overline{w},u^{\sigma},v^{\sigma},w^{\sigma},u^{\dagger},v^{\dagger},w^{\dagger}\in R[x]$
@@
$h$ $=(I_{n}+V_{n-1}+x)u-v-w$
$=xu+I_{n}u_{V_{n}}+V_{n-1}u_{I_{n-1}}$
we only need to show that $u=|\mathcal{L}_{1}|$ is subtraction-free. This follows from the tridiagonal structure of $|\mathcal{L}_{1}|$ (and Lemma 2.1), as one can immediately see that the only negative terms appearing $|\mathcal{L}_{1}|$ are of the form $I_{j}V_{j}|\mathcal{L}_{n-j+1}|$ and those are cancelled by the expressions $(I_{j}+V_{j-1}+x)(I_{j+1}+V_{j}+x)|\mathcal{L}_{n-j+1}|$. ∎
@@
The last Lemma is the main reason for our choice of normalization of $\mathcal{C}$, as subtraction-free polynomials can be nicely tropicalized. This will be used in appendix C.
@@
$|{}^{1}\mathcal{L}_{1}||\mathcal{L}_{2}|-|\mathcal{L}_{1}||{}^{1}\mathcal{L}_{2}|=I_{1}V_{1}\cdots I_{n-2}V_{n-2}.$
@@
$|{}^{2}\mathcal{L}_{1}||{}^{1}\mathcal{L}_{2}|-|{}^{1}\mathcal{L}_{1}||{}^{2}\mathcal{L}_{2}|=I_{2}V_{2}\cdots I_{n-2}V_{n-2}.$
@@
$|{}^{n-2}\mathcal{L}_{1}||{}^{n-3}\mathcal{L}_{2}|-|{}^{n-3}\mathcal{L}_{1}||{}^{n-2}\mathcal{L}_{2}|=I_{n-2}V_{n-2}.$
@@
$I_{n}V_{n}|{}^{1}\mathcal{L}_{2}|=(I_{n}+V_{n-1}+x)v^{\sigma^{-1}}-u^{\sigma^{-1}}$
@@
This follows from $|{}^{1}\mathcal{L}_{2}|=|{}^{2}\mathcal{L}^{\sigma^{-1}}_{1}|$ and Lemma 2.1. ∎
@@
$I_{n}u-v$ $=I_{n}u_{V_{n}}$
$=I_{n}\overline{u}-\overline{w},$
$(V_{n-1}+x)u-w$ $=(V_{n}+x)\overline{u}-\overline{v}.$
@@
$u$ $=(I_{1}+V_{n}+x)|{}^{1}\mathcal{L}_{1}|-I_{1}V_{1}|{}^{2}\mathcal{L}_{1}|,$
$\overline{u}$ $=(I_{1}+V_{1}+x)|{}^{1}\overline{\mathcal{L}}_{1}|-I_{2}V_{1}|{}^{2}\overline{\mathcal{L}}_{1}|$
and Lemma 2.4. The second assertion follows from the first and $h=\overline{h}.$
@@
$u_{n}$ $=(u_{n})_{V_{n}}+V_{n}u_{n-1}^{\sigma_{n-1}},$
$v_{n}$ $=I_{n}V_{n}u_{n-1}^{\sigma_{n-1}},$
$w_{n}$ $=I_{n-1}V_{n-1}u_{n-1}^{\sigma_{n-1,n}}.$
@@
We start with the second relation. This follows from
$|{}^{1}\mathcal{L}_{1}|=|\mathcal{L}^{\sigma_{n}}_{2}|=|(\mathcal{L}_{n-1}^{\sigma_{n-1}})_{1}|.$
@@
$u_{n}-(u_{n})_{V_{n}}=u_{n}-u_{n}+V_{n}|{}^{1}\mathcal{L}_{1}|$
and the second relation. The other relations follow similarly. ∎
Remark 2.8.
In particular, we have
$(I_{n}+V_{n-1}+x)\,u_{n}=h_{n}+I_{n}V_{n}u_{n-1}^{\sigma_{n-1}}+I_{n-1}V_{n-1}u_{n-1}^{\sigma_{n-1,n}}.$
@@
$(\overline{u}_{n})_{V_{n-1}}=(u_{n})_{V_{n}}.$
2.6. Mumford’s algebraic description of $\mathrm{Jac(\mathcal{C})}$